# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ B.1 Example 1
Another property to distinguish between RD and TD is the magnitude of the normal magnetic field component $B_{n}$. The divergence-free condition implies that $B_{n}$ must be constant in case of planar geometry. So, MVA minimum variance direction should represent normal to the magnetic sheet, and $B_{3}=B_{n}=$ constant. Next, tangential discontinuities have $B_{n}=0$, but in observations $B_{3}\simeq 70$ nT. These results are obtained in the known limits of MVA since the MVA estimation of the normal to the sheet can differ from normal estimated from multi-spacecraft methods (Horbury et al., 2001; Knetter et al., 2004).
So, to summarize, starting from the largest observed scales and up to the end of the inertial range, we observe a current sheet that can be interpreted in terms of a tangential discontinuity (TD). At ion and sub-ion scale substructures are embedded in this discontinuity. Ion scale structure resembles the dipole Alfvén vortex model (see Section 5 and Figure 11 column (b)). Sub-ion scale structure might represent a compressible vortex (Jovanović et al., 2020). A sketch describing this event is given in Figure 10.
B.2 Example 2
The second example is shown in Figure 12 in the same format as the first event in Figure 11. The central time of the event is 01:19:20 UT. In panel (a) the raw magnetic field is shown in RTN reference frame.
Panel (e) shows the polarization $B_{T}(B_{N})$; out-of-plane $B_{R}$ does not change sign (this structure is not a switchback). The magnetic field deflects twice within the timescale of $\sim 80$ s. Magnetic field on the left side of $t_{0}$ is different from the right hand side: it rotates by $\Delta\zeta=15^{\circ}$ (see Equation (A3)). This can be due to a weak ($|\Delta\mathbf{B}|/B_{0}=(\mathbf{B}_{t=-100s}-\mathbf{B}_{t=100s})/B_{0}\simeq 0.3$) rotational current sheet, since the ratio of velocity and magnetic field jumps satisfy the Walen relation |$\Delta\mathbf{V}|/V_{A}=1.03\cdot|\Delta\mathbf{B}|/B_{0}$.
Magnetic fluctuations at the MHD scales $\delta B_{MHD}$ are shown in Figure 12 (b) in the MVA reference frame. The amplitude of the structure $\delta B_{1}=60$ nT (i.e., from peak to peak $\Delta B/B_{0}\sim 1.2$) well exceeds the level of incoherent signal $222\sigma_{MHD}=$ 22 nT. The direction of the maximum eigenvector $\mathbf{e}_{1}$ is well-distinguished from intermediate ($\mathbf{e}_{2}$) and minimum ($\mathbf{e}_{3}$) directions since $\lambda_{2}/\lambda_{1}=0.06$, and it is perpendicular to the background magnetic field $\theta_{B_{0},1}=88^{\circ}$.
The velocity and magnetic field fluctuations are well-correlated $\delta B/B_{0}=\xi\delta V/V_{A}$, see panels (i-k). The proportionality coefficient is $\xi=0.86$ for magnetic fluctuations at MHD scales and $\xi=0.81$ in the raw data with the mean value subtracted for the same time interval. The black lines in panels (i-k) show the fluctuations of the monopole Alfvén vortex model, along the central crossing. The observed time profiles of fluctuations agree well with the Alfven vortex model. The magnetic field vector before and after the vortex crossing is oriented differently: $|\mathbf{B}(t_{0}-70~{}\text{s})-\mathbf{B}(t_{0}+70~{}\text{s})|/B_{0}=(0.4,0.03,0.03)$ (in the MVA reference frame). Thus, the vortex is embedded in a rather weak current sheet.
In the plasma rest frame this vortex propages with a negligeble speed of $\simeq 3$ km/s (see the main text of the paper), so the Taylor hypothesis can be used to estimate its spatial scale. PSP trajectory crosses the structure in a plane nearly perpendicular to $B_{0}$ ($\theta_{B_{0},V}=71^{\circ}$). The diameter of the vortex can be estimated as $d=\theta_{B_{0},V}\cdot V\Delta t\sim 2.4\cdot 10^{4}$ km. Given that the mean ion inertial length is $d_{i}\simeq 11$ km and the mean ion Larmor radius is $\rho_{i}=5$ km, the vortex diameter $d\simeq 2\cdot 10^{3}d_{i}=5\cdot 10^{3}\rho_{i}$. The variation of the magnetic field modulus is negligible $\delta|B|/B_{0}=0.03$ as well as the variations of $N_{e}$, $T_{e}$ and $T_{p}$, see panels (l-n). The change of $|V|$ (grey line in panel (l)) is due to the superposition of the velocity fluctuation of the Alfvén vortex on the bulk solar wind speed.
Figure 12(c) shows the ion scale magnetic fluctuations $\delta B_{ion}$ located in the center of the MHD scale Alfvén vortex. The maximum amplitude of the fluctuation: $\max(|\delta B_{1}|)=24$ nT, as well as two secondary peaks on the left and right sides exceed well the incoherent threshold ±$\pm 2\sigma_{ion-scale}$ shown in grey. The polarization is elliptical (panel (g)), and the maximum variance is perpendicular to the local field direction. These observed properties are in agreement with an Alfvén vortex crossing with a finite impact distance from its center. The vortex was observed during $\Delta t=1.5$ s. The diameter of the vortex is $d=V\Delta t\sin(\Theta_{BV})=440$ km $=40~{}d_{i}=90~{}\rho_{i}$, where $\Delta t=1.5$ s is the duration of the crossing, $V=330$ km/s, $\Theta_{BV}=70^{\circ}$, $d_{i}=11$ km, $\rho_{i}=5$ km.
[…]
@@ B.2 Example 2
In summary, in the example of Figure 12, in raw data we observe a weak current sheet with the thickness of the high amplitude MHD scale structure embedded in it. This current sheet is Alfvénic in nature that is the property of a rotational discontinuity. We can interpret the MHD structure embedded in the weak rotational current sheet as a monopole Alfvén vortex crossed close to its center. Within this monopole Alfvén vortex, we observe smaller scale vortices at ion and sub-ion scales. The sketch of the Example 2 is presented in the Figure 10 (b).
@@ B.3 Example 3
Figure 13 shows an example observed around $t_{0}=$ 2:19:38 UT. In panel (a) the center of the current sheet is observed at $t-t_{0}=20$ s, when the magnetic field rotates by the angle $\Delta\zeta=110^{\circ}$. $B_{R}$ changes sign across the sheet, so the sheet forms the boundary of a switchback, similarly to observations in (Krasnoselskikh et al., 2020).
@@
In Figure 13(b), MHD scale fluctuations in MVA reference frame are shown. The amplitude of fluctuations, associated with the discontinuity, exceed the level of the incoherent signal (see grey horizontal band). Panels (e-f) show the corresponding polarizations.
@@
The fluctuations are nearly Alfvénic in the vicinity of the discontinuity: $(\mathbf{B}-\langle\mathbf{B}\rangle)_{1}/B_{0}=1.25\cdot(\mathbf{V}-\langle\mathbf{V}\rangle)_{1}/V_{A}$, when $t-t_{0}\in(0,40)$ s, see Figure 13(i-k). But further away from the discontinuity $\Delta B/B_{0}$ and $\Delta V/V_{A}$ have different amplitudes: $\Delta B_{1}/B_{0}\approx 2.2\cdot\Delta V_{1}/V_{A}$, where $\Delta B_{1}=|B_{1}(t_{0}-100\text{\,s})-B_{1}(t_{0}+100\text{\,s})|$ and $\Delta V_{1}=|V_{1}(t_{0}-100\text{\,s})-V_{1}(t_{0}+100\text{\,s})|$. The magnetic field modulus decreases from 105 nT at the boundaries to 90 nT in the center (panel (l)). The duration of this magnetic cavity is $\Delta t\simeq 100$ s, which corresponds to the scale of $l=\Delta t\cdot V=3.5\cdot 10^{4}$ km $=3\cdot 10^{3}d_{i}=5\cdot 10^{3}\rho_{i}$. The density $N_{e}$, Figure 13 (m), weakly increases across the discontinuity. The proton temperature $T_{p}$, Figure 13 (n), is higher on the left side of the discontinuity than on the right, it has a local maximum around the discontinuity center in contrast with a nearly uniform $T_{e}$. There is an anti-correlation between the norm of the field and plasma density $N_{e}$, indicating a possible pressure balance. The polarization of magnetic fluctuations is arch-like, that is typical for rotational discontinuities (Tsurutani et al., 1996; Sonnerup et al., 2010; Haaland et al., 2012; Paschmann et al., 2013).
@@
Magnetic fluctuations at ion scales are shown in Figure 13 (c). The maximum and intermediate magnetic fluctuations ($\delta B_{1}$ and $\delta B_{2}$) are transverse and have nearly the same amplitude, the polarisation is close to elliptic, Figure 13 (g). The described properties (i.e. localised transverse fluctuations with nearly elliptic polarisation) are consistent with the off-center monopole Alfvén vortex crossing, see Figure 6 (a). The minimum MVA eigenvector $\textbf{e}_{3}$ is well-defined, because $\lambda_{3}/\lambda_{2}=0.06$ is small. From the model crossings (see Section 5.1.1) we know that if the spacecraft is crossing the monopole Alfven vortex, $\textbf{e}_{3}$ is a good approximation for the axis of the vortex. We can conclude that the axis of the vortex is nearly parallel to the background magnetic field, since $\theta_{B_{0},3}=8^{\circ}$.
@@
The sub-ion scale structure, Figure 13(d), is well localised in time and has a significantly compressible component, $\delta|B|\sim 0.5\,\delta B_{1}$. This high compressibility is a typical properties of structures at these scales. Such localized compressible magnetic fluctuations at sub-ion scales can be interpreted as the compressible Alfvén vortex (Jovanović et al., 2020).
@@ B.4 Example 4
Figure 14 shows the details of the fourth example. In panel (a) the radial magnetic field $B_{R}$ is positive in two time intervals: $t-t_{0}\in(-30,0)$ s and $t-t_{0}\in(25,80)$ s. These are two neighbouring switchbacks. The crossing is nearly perpendicular to the mean background magnetic field $\mathbf{B_{0}}$ ($\theta_{BV}=71^{\circ}$), since $\mathbf{B_{0}}=\langle\mathbf{B}\rangle_{t-t_{0}\in(-100,100)~{}s}=(11,84,-14)$ nT is directed mostly along the tangential direction in the RTN reference frame and the solar wind bulk speed is $\mathbf{V}=\langle\mathbf{V}\rangle_{t-t_{0}\in(-100,100)~{}s}=(400,80,0)$ km/s.
Figure 14(b) shows the MHD scale bandpass filtered magnetic fluctuations in MVA reference frame. We observe sharp discontinuities at the boundaries of both neighbouring switchbacks. The magnetic field and velocity fluctuations are well-correlated, the normalised amplitudes of $\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ are equal, see panels (i-k), so the fluctuations are Alfvénic. The density $N_{e}$ is constant across the structure, see panel (m). The proton temperature $T_{p}$ is increased by $\sim 30$% in both neighbouring switchbacks compared to the value outside, see panel (n).
Similar to the previous examples, the embedded structures are observed in the vicinity of the central time of the event $t-t_{0}=0$ s (at the right boundary of the first switchback). Fluctuations of ion scales are shown in the panel (c). The polarization is close to linear, see panel (g). The maximum variance direction is perpendicular to the local background magnetic field. The time profile of the corresponding component $\delta B_{1}$ is consistent with the monopole Alfvén vortex crossing through the center.
Figure 14(d) shows the sub-ion fluctuations. The profile of the maximum variance component $\delta B_{1}$ has a peak. The intermediate component $\delta B_{2}$ coincide with $\delta|B|$, so the structure is compressible, with $\delta|B|/\max{|\delta B_{1}|}\sim 0.3$. It can be interpreted in terms of compressible ion scale vortex (Jovanović et al., 2020).
@@ B.5 Summary of detected structures
We collected large statistics of coherent structures (Figure 4). At MHD scales some of these events represent isolated current sheets such as tangential and rotational current sheets with two examples shown in Section B.1 and B.3 respectively. However, we found that current sheets are rare, most of the events can be interpreted as Alfvén vortices. The example in Section B.2 (Figure 12) is interpreted as the crossing of a monopole vortex along its center (embedded in a weak and large scale rotationnal discontinuity). From the visual examination of $\sim 200$ events, we find that the embedded structures at ion and sub-ion scales are mostly Alfvén vortices, independently on the existence of a CS at large scales. From the examples shown here, in case of CS at large scales, the sub-ion vortices are compressible and in the case of the large scale Alfvén vortex, the small scale vortex is incompressible.
@@ Appendix C Necessity of band-pass filtering
At ion and sub-ion scales the amplitude of embedded substructures are much smaller than the amplitude of MHD-scale fluctuations and the the background magnetic field. Therefore, filtration allows to remove the quasi-constant background and to analyze specifically the fluctuations associated with substructures. However, the filtration may introduce an ambiguity in the interpretation of the signal. Let us consider the thin (i.e. $\ell\sim d_{i}$) current sheet, so that the crossing duration is $\Delta t=0.125$ s. Figure 15 shows the tangential magnetic field component of the current sheet (blue) and the result of the filtration (orange).
In Figure 15(a) the filter frequency window corresponds to the MHD inertial range, as defined in Section 3 in Equation  (6). Since the thickness of the sheet is small, $\Delta t/\text{min}(\tau_{MHD})=0.1$, the filtered signal has a steep jump of the same amplitude as the amplitude of the initial signal. However, unlike the initial signal, the filtered signal tends to 0 at the scale $|t|>\text{max}(\tau_{MHD})/2$ away from the discontinuity. Two low amplitude local extreme appear at $t=\pm 50$ s.
Figure 15(b) shows the result of the filtration at ion scales. The magnetic field changes sign smoothly. If the thickness of an intense coherent structure is smaller than the minimum timescale of the MHD range, and if the classification method is based on the shape of the most intense filtered magnetic field component, then the CS filtration remnant at ion scales can be miss-classified as an embedded monopole Alfvén vortex crossed through its center.
In conclusion, filtering is necessary for the study of ion and subion scale structures because such structures have small amplitude compared to $B_{0}$, and they are poorly distinguishable in raw data. However, filtering can significantly change the waveform, which complicates direct comparison of structures with models. For example, the current sheet does not look like a step function after filtering. However, signal filtering has little effect on the polarisation of fluctuations. For example, if the spacecraft is crossing a tangential discontinuity (see Eq. A4) with only one component of the magnetic field changed, i.e. with linear polarisation, then after filtering, the shape of the signal will change, but the polarisation will remain linear. If the spacecraft is crossing an Alfvén vortex and the measured polarisation is close to elliptical in raw data, then filtering will remove the quasi-constant background magnetic field, but the polarisation will remain elliptical. The polarisation is convenient to show the amplitude anisotropy of a 2D vector. But in general case the magnetic fluctuations is a 3D vector. So the amplitude anisotropy can be characterised by MVA eigenvalue relations.

### v4

@@ B.2 Example 2
In summary, in the example of Figure 12, in raw data we observe a weak current sheet with the thickness of the high amplitude MHD scale structure embedded in it. This current sheet is Alfvénic in nature that is the property of a rotational discontinuity. We can interpret the MHD structure embedded in the weak rotational current sheet as a monopole Alfvén vortex crossed close to its center. Within this monopole Alfvén vortex, we observe smaller vortices at ion and sub-ion scales. Figure 10 (b) shows the sketch of the Example 2.
@@ B.3 Example 3
Figure 13 shows an example observed around $t_{0}=$ 2:19:38 UT. In panel (a), the center of the current sheet is observed at $t-t_{0}=20$ s, when the magnetic field rotates by the angle $\Delta\zeta=110^{\circ}$. $B_{R}$ changes sign across the sheet, so the sheet forms the boundary of a switchback, similarly to observations in (Krasnoselskikh et al., 2020).
@@
Figure 13(b) shows MHD scale fluctuations in the MVA reference frame. The amplitude of fluctuations associated with the discontinuity exceeds the level of the incoherent signal (see grey horizontal band). Panels (e-f) show the corresponding polarizations.
@@
The fluctuations are nearly Alfvénic in the vicinity of the discontinuity: $(\mathbf{B}-\langle\mathbf{B}\rangle)_{1}/B_{0}=1.25\cdot(\mathbf{V}-\langle\mathbf{V}\rangle)_{1}/V_{A}$, when $t-t_{0}\in(0,40)$ s, see Figure 13(i-k). But further away from the discontinuity $\Delta B/B_{0}$ and $\Delta V/V_{A}$ have different amplitudes: $\Delta B_{1}/B_{0}\approx 2.2\cdot\Delta V_{1}/V_{A}$, where $\Delta B_{1}=|B_{1}(t_{0}-100\text{\,s})-B_{1}(t_{0}+100\text{\,s})|$ and $\Delta V_{1}=|V_{1}(t_{0}-100\text{\,s})-V_{1}(t_{0}+100\text{\,s})|$. The magnetic field modulus decreases from 105 nT at the boundaries to 90 nT in the center (panel (l)). The duration of this magnetic cavity is $\Delta t\simeq 100$ s, which corresponds to the scale of $l=\Delta t\cdot V=3.5\cdot 10^{4}$ km $=3\cdot 10^{3}d_{i}=5\cdot 10^{3}\rho_{i}$. The density $N_{e}$, Figure 13 (m), weakly increases across the discontinuity. The proton temperature $T_{p}$, Figure 13 (n), is higher on the left side of the discontinuity than on the right, it has a local maximum around the discontinuity center in contrast with a nearly uniform $T_{e}$. There is an anti-correlation between the magnetic field modulus and plasma density $N_{e}$, indicating a possible pressure balance. The polarization of magnetic fluctuations is arch-like, which is typical for rotational discontinuities (Tsurutani et al., 1996; Sonnerup et al., 2010; Haaland et al., 2012; Paschmann et al., 2013).
@@
Magnetic fluctuations at ion scales are shown in Figure 13 (c). The maximum and intermediate magnetic fluctuations ($\delta B_{1}$ and $\delta B_{2}$) are transverse and have nearly the same amplitude, the polarization is close to elliptic, Figure 13 (g). The described properties (i.e. localized transverse fluctuations with nearly elliptic polarization) are consistent with the off-center monopole Alfvén vortex crossing, see Figure 6 (a). The minimum MVA eigenvector $\textbf{e}_{3}$ is well-defined, because $\lambda_{3}/\lambda_{2}=0.06$ is small. From the model crossings (see Section 5.1.1) we know that if the spacecraft is crossing the monopole Alfven vortex, $\textbf{e}_{3}$ is a good approximation for the axis of the vortex. We can conclude that the axis of the vortex is nearly parallel to the background magnetic field, since $\theta_{B_{0},3}=8^{\circ}$.
@@
The sub-ion scale structure, Figure 13(d), is well localized in time and has a significantly compressible component, $\delta|B|\sim 0.5\,\delta B_{1}$. This high compressibility is a typical property of structures at these scales. Such localized compressible magnetic fluctuations at sub-ion scales can be interpreted as the compressible Alfvén vortex (Jovanović et al., 2020).
@@ B.4 Example 4
Figure 14 shows the details of the fourth example. In panel (a) the radial magnetic field $B_{R}$ is positive in two time intervals: $t-t_{0}\in(-30,0)$ s and $t-t_{0}\in(25,80)$ s. These are two neighboring switchbacks. The crossing is nearly perpendicular to the mean background magnetic field $\mathbf{B_{0}}$ ($\theta_{BV}=71^{\circ}$), since $\mathbf{B_{0}}=\langle\mathbf{B}\rangle_{t-t_{0}\in(-100,100)~{}s}=(11,84,-14)$ nT is directed mainly along the tangential direction in the RTN reference frame and the solar wind bulk speed is $\mathbf{V}=\langle\mathbf{V}\rangle_{t-t_{0}\in(-100,100)~{}s}=(400,80,0)$ km/s.
Figure 14(b) shows the MHD scale bandpass filtered magnetic fluctuations in the MVA reference frame. We observe sharp discontinuities at the boundaries of both neighboring switchbacks. The magnetic field and velocity fluctuations are well-correlated, the normalized amplitudes of $\mathbf{B}-\langle\mathbf{B}\rangle)/B_{0}$ and $\mathbf{V}-\langle\mathbf{V}\rangle)/V_{A}$ are equal, see panels (i-k), so the fluctuations are Alfvénic. The density $N_{e}$ is constant across the structure, see panel (m). The proton temperature $T_{p}$ is increased by $\sim 30$% in both neighboring switchbacks compared to the value outside, see panel (n).
Similar to the previous examples, the embedded structures are observed near the central time of the event $t-t_{0}=0$ s (at the right boundary of the first switchback). Fluctuations of ion scales are shown in the panel (c). The polarization is close to linear, see panel (g). The maximum variance direction is perpendicular to the local background magnetic field. The time profile of the corresponding component $\delta B_{1}$ is consistent with the monopole Alfvén vortex crossing through the center.
Figure 14(d) shows the sub-ion fluctuations. The profile of the maximum variance component $\delta B_{1}$ has a peak. The intermediate component $\delta B_{2}$ coincide with $\delta|B|$, so the structure is compressible, with $\delta|B|/\max{|\delta B_{1}|}\sim 0.3$. We interpret this structure as a compressible ion scale vortex (Jovanović et al., 2020).
@@ B.5 Summary of detected structures
We collected large statistics of coherent structures (Figure 4). Some of these events at MHD scales represent isolated current sheets such as tangential and rotational current sheets, see two examples shown in Section B.1 and B.3 respectively. However, we found that current sheets are rare, most of the events can be interpreted as Alfvén vortices. The example in Section B.2 (Figure 12) is interpreted as the crossing of a monopole vortex along its center (embedded in a weak and large scale rotational discontinuity). From the visual examination of $\sim 200$ events, we find that the embedded structures at ion and sub-ion scales are mostly Alfvén vortices, independently on the existence of a CS at large scales. From the examples shown here, in the case of CS at large scales, the sub-ion vortices are compressible and in the case of the large scale Alfvén vortex, the small scale vortex is incompressible.
@@ Appendix C Necessity of band-pass filtering
At ion and sub-ion scales the amplitude of embedded substructures is much smaller than the amplitude of MHD-scale fluctuations and the background magnetic field. Therefore, filtration allows to remove the quasi-constant background, which facilitates the analysis of the fluctuations associated with substructures. However, the filtration may introduce an ambiguity in the interpretation of the signal. Let us consider the thin (i.e. $\ell\sim d_{i}$) current sheet, so that the crossing duration is $\Delta t=0.125$ s. Figure 15 shows the tangential magnetic field component of the current sheet (blue) and the result of the filtration (orange).
In Figure 15(a) the filter frequency window corresponds to the MHD inertial range, as defined in Section 3 in Equation  (6). Since the thickness of the sheet is small, $\Delta t/\text{min}(\tau_{MHD})=0.1$, the filtered signal has a steep jump of the same amplitude as the amplitude of the initial signal. However, unlike the initial signal, the filtered signal tends to 0 at the scale $|t|>\text{max}(\tau_{MHD})/2$ away from the discontinuity. Two low amplitude local extremes appear at $t=\pm 50$ s.
Figure 15(b) shows the result of the filtration at ion scales. The magnetic field changes the sign smoothly. If the thickness of an intense coherent structure is smaller than the minimum timescale of the MHD range, and if the classification method is based on the shape of the most intense filtered magnetic field component, then the CS filtration remnant at ion scales can be miss-classified as an embedded monopole Alfvén vortex crossed through its center.
In conclusion, filtering is necessary for the study of ion and subion scale structures because such structures have small amplitude compared to $B_{0}$, and they are poorly distinguishable in raw data. However, filtering can significantly change the waveform, which complicates the direct comparison of structures with models. For example, the current sheet does not look like a step function after filtering. However, signal filtering has little effect on the polarization of fluctuations. For example, if the spacecraft is crossing a tangential discontinuity (see Eq. A4) with only one component of the magnetic field changed, i.e. with linear polarization, then after filtering, the shape of the signal will change, but the polarization will remain linear. Suppose the spacecraft is crossing an Alfvén vortex and the measured polarization is close to elliptical in raw data. In that case, filtering will remove the quasi-constant background magnetic field, but the polarization will remain elliptical. The polarization is convenient to show the amplitude anisotropy of a 2D vector. But in general case the magnetic fluctuations are a 3D vector. So, the amplitude anisotropy can be characterized by MVA eigenvalue relations.